### v2

# Epidemiology of Large Language Models: A Benchmark for Observational Distribution Knowledge

Abstract
Artificial intelligence (AI) systems hold great promise for advancing various scientific disciplines, and are increasingly used in real-world applications. Despite their remarkable progress, further capabilities are expected in order to achieve more general types of intelligence. A critical distinction in this context is between factual knowledge, which can be evaluated against true or false answers (e.g., “what is the capital of England?”), and probabilistic knowledge, which reflects probabilistic properties of the real world (e.g., “what is the sex of a computer science graduate in the US?”). Much of previous work on evaluating large language models (LLMs) focuses on factual knowledge, while in this paper, our goal is to build a benchmark for understanding the capabilities of LLMs in terms of knowledge of probability distributions describing the real world. Given that LLMs are trained on vast amounts of text, it may be plausible that they internalize aspects of these distributions. Indeed, this idea has gained traction, with LLMs being touted as powerful and universal approximators of real-world distributions. At the same time, classical results in statistics, known under the term curse of dimensionality, highlight fundamental challenges in learning distributions in high dimensions, challenging the notion of universal distributional learning. In this work, we develop the first benchmark to directly test this hypothesis, evaluating whether LLMs have access to empirical distributions describing real-world populations across domains such as economics, health, education, and social behavior. Our results demonstrate that LLMs perform poorly overall, and do not seem to internalize real-world statistics naturally. This finding also has important implications that can be interpreted in the context of Pearl’s Causal Hierarchy (PCH). Our benchmark demonstrates that language models do not contain knowledge on observational distributions (Layer 1 of the PCH), and thus the Causal Hierarchy Theorem implies that interventional (Layer 2) and counterfactual (Layer 3) knowledge of these models is also limited.
footnotetext: Project page available at https://llm-observatory.org/.
Keywords: generative artificial intelligence, large language models, probabilistic knowledge
1 Introduction
Artificial intelligence (AI) systems hold great potential to accelerate many scientific disciplines including healthcare (jiang2017artificial; shaheen2021applications), education (holmes2022state), and economics (bickley2022artificial), just to name a few. Automated systems based on AI are also increasingly used in a wide variety of real-world settings. Yet, despite rapid progress and broad adoption of these systems, many core capabilities required for robust and reliable AI still need to be developed further, on the path to reaching artificial human intelligence (bubeck2023sparks; kaplan2020scaling; hendrycks2025definition). One of the hallmark features expected from AI systems is the ability to reason probabilistically about the world, and not only to mimic human intuition and language. Such capabilities are needed for systems capable of assisting scientific discovery, advancing our understanding of complex phenomena, and help societal decision-making.
In this context, it is helpful to distinguish between two types of knowledge embedded in AI systems: factual knowledge, which can be evaluated based on true or false answers based on questions for which there is an agreed upon answer (e.g., “what is the capital of England?”); and probabilistic knowledge, which reflects the inherent uncertainties of the world we operate in (e.g., question such as “what is the sex of a computer science graduate in the US?”). Much of the current evaluations of large language models (LLMs), currently the prevailing paradigm in AI research (bommasani2021opportunities), has focused on factual knowledge, i.e., question answering, fact recall, etc. At the same, the degree to which LLMs encode probabilistic knowledge about the real world remains relatively underexplored.
In this work, our focus is on observational probabilistic knowledge, where the term observational refers to the first layer of Pearl’s Causal Hierarchy (PCH, for short) (pearl:2k; bareinboim2020on). The PCH distinguishes different types of probabilistic knowledge, corresponding to different cognitive capabilities of (AI) systems: Layer 1 of observation (seeing), which includes reasoning about observed events and correlations between variables; Layer 2 of intervention (doing), which includes inferences about systems in which causal interventions are performed; and Layer 3 of counterfactual reasoning (imagining, see Fig. 2 left), which allows inference for hypothesized events for which the required pre-conditions may not have materialized in reality. Important results from previous work demonstrate that, in absence of appropriate assumptions, moving across different layers of PCH is provably impossible. The result known as the Causal Hierarchy Theorem (CHT, (bareinboim2020on)) states that observational knowledge (Layer 1) about a system underdetermines its interventional (Layer 2) and counterfactual (Layer 3) behavior. With causal modeling and assumptions, inferences across layers may become feasible. A classical example is causal modeling that uses observational data, combined with structural assumptions, to infer interventional or counterfactual distributions (pearl:2k; bareinboim2016causal). Therefore, if one can show that LLMs do not have access to observational knowledge (Layer 1), based on the CHT, one may also be skeptical of these models’ capabilities for reliable statements about interventions or counterfactuals.
Given the abundance of observational data, it is plausible that LLMs, trained on vast corpora of text data, have internalized aspects of real-world observational distributions. Indeed, strong claims in this debate have emerged, with OpenAI’s CEO Sam Altman suggesting that “humanity has found a universal way to approximate distributions,” (altman2024intelligence) implying that large models serve as powerful approximators of the world, and may have access to any kind of Layer 1 knowledge.
While LLMs have been tremendously successful at various tasks, some caution is nonetheless warranted. Classical results in statistics highlight fundamental limitations of learning distributions in higher dimensions: the famous curse of dimensionality results demonstrate theoretically that learning distributions becomes exponentially harder (in terms of required samples) as dimensionality grows. Seminal work by stone1982optimal establishes that estimation rates degrade sharply with dimensionality, with the optimal rate of convergence for an unknown differentiable regression function $f$ being $\mathcal{O}(n^{-\frac{1}{2+d}})$ in any $L^{p}$ norm, where $d$ is the dimension of the input of $f$. Such results challenge the notion of universal distributional learning in higher dimensions, in stark contrast with the notion of LLMs working as universal approximators.
Figure 1: Benchmark leaderboard.
This tension raises a natural question: should one believe the optimism of universal approximation put forward by Sam Altman, or the caution grounded in statistical theory, such as the results of Charles Stone? In this paper, we attempt to shed light on this debate and examine different aspects of LLMs’ capabilities in approximating distributions. We introduce a benchmark specifically designed to assess whether LLMs have access to observational distributions describing real-world populations, across various domains including economics, health, education, crime, and social behavior. This is why our approach can be termed as epidemiology, from Greek epi (fall upon) + demos (population) + logos (knowledge) – the knowledge of what falls upon populations.
1.1 Contributions
Our contribution in this work is to construct a benchmark for evaluating the capabilities of large language models in terms of access to knowledge about observational distributions in the real-world.
  1. (i)
    We curate a total of 10 datasets, ranging from healthcare, health behavior, and education, to labor, consumer spending, and crime statistics. For each large scale dataset, which describes the population level statistics in the United States, we extract a set of questions to test whether a language model has knowledge about the population encoded in the respective dataset.
  2. (ii)
    We design a modular evaluation framework that facilitates easy benchmarking of additional models, making it straightforward to add datasets, questions, or models to our benchmark.
A preview of our results is shown in Fig. 1, indicating that over a range of models, the performance of the current generation of LLMs in terms of knowledge of real-world distributions is quite poor. Through an example, we emphasize why this type of knowledge may be important for the end user:
Example 1 (Diabetes Rates)
A user with persistent fatigue is unsure whether their symptoms may be indicative of diabetes. They provide an LLM with basic demographic and lifestyle information, such as age, sex, BMI, smoking status, and ask: “Given my age (45), female sex, BMI (29), non-smoker status, what is the baseline probability that I have diabetes?”
Answering the query above reliably requires access to the observational distribution governing disease risk in the population, and our later investigation demonstrates that current models do not possess this knowledge. The illustrative example is just one instance of a query where observational knowledge is essential, and similar type of knowledge is relevant across many other domains investigated in this manuscript.
Figure 2: Overview figure for our benchmark.
Related Literature.
A growing body of literature investigates aspects of knowledge and generalization abilities of LLMs. The two works most related to ours are that of santurkar2023whose, which examines the responses on language models on Pew Research Center’s American Trends Panel (pew2024atp), in order to investigate if LLMs exhibit responses of a specific demographic group. Another related work is that of dominguez2024questioning, that examines the responses of language models on the American Community Survey (ACS) (acs2024census), and compares them to the real responses of individual. Our work takes a broader view, and attempts to build a benchmark for systematically analyzing the knowledge of models on observational distributions (Layer 1 of the PCH), across a wide range of datasets, and distributions over variables with different dimensions.
Other works investigate the ability of models to perform correct probabilistic inferences, given information in forms of probability tables or probability conditionals (he2023solving; jin2023cladder; nafar2024probabilistic). This line of work is distinct from ours, since our benchmark aims to evaluate the availability of the true probabilities to the model, rather than the validity of inferences over these probabilities. We also mention the work of zhao2023felm, aimed at testing factual correctness of model responses on various domains, including knowledge about the world, math, or reasoning. In parallel, other lines of work use the term faithfulness while evaluating language models, usually referring to faithfulness of model responses to the the actual source material where the information is available. For instance, (zhou2023context) investigates whether responses of models are affected by knowledge conflict, and whether they exhibit appropriate response abstentions. This work is focused on testing factual knowledge, as opposed to descriptive knowledge investigated by our benchmark. Another body of work looks at faithfulness of model explanations, that is, whether the explanation cited by the model is in fact compatible with the model’s actual reasoning steps (agarwal2024faithfulness; matton2025walk), differing from the approach and the goal of our work.
Relationship to Causal Inference.
While not directly focused on causal knowledge, our investigation still bears important implications for causal inference, which we now elucidate. Following the standard approach in the literature (pearl:2k), a structural causal model (SCM) is a tuple $\mathcal{M}:=\langle V,U,\mathcal{F},P(u)\rangle$, where $V$, $U$ are sets of endogenous (observable) and exogenous (latent) variables, respectively, $\mathcal{F}$ is a set of functions $f_{V_{i}}$, one for each $V_{i}\in V$, where $V_{i}\leftarrow f_{V_{i}}(\mathrm{pa}(V_{i}),U_{V_{i}})$ for some $\mathrm{pa}(V_{i})\subseteq V$ and $U_{V_{i}}\subseteq U$. The assignment mechanisms $\mathcal{F}$ determine how each variable attains its value, and the set $\mathrm{pa}(V_{i})$ is called the parent set of $V_{i}$. Together with the probability distribution $P(u)$ over the exogenous variables $U$, the SCM specifies the entire behavior of the underlying phenomenon, meaning that it fully specifies all observational, interventional, and counterfactual distributions (bareinboim2020on).
The result known as Causal Hierarchy Theorem (bareinboim2020on) shows formally that access to observational distributions is a fundamental step towards performing interventional (Layer 2) or counterfactual (Layer 3) inference:
Proposition 1 (Causal Hierarchy Theorem (bareinboim2020on))
Let $\mathcal{M}$ be an SCM, $P(V)$ be its observational distribution, and $\mathcal{A}$ a set of causal assumptions encoded in the form of a causal diagram (bareinboim2020on) or ignorability statements (rubin1974estimating; pearl:2k). Then,
  1. (a)
    in absence of $\mathcal{A}$, $P(V)$ underdetermines interventional and counterfactual distributions,
  2. (b)
    in absence of $P(V)$, $\mathcal{A}$ underdetermines interventional and counterfactual distributions.
The first part of the above proposition is the one commonly considered in the causal inference literature – in short, it states that, in absence of causal assumptions, it is generally impossible to provide any guarantees for inference over interventional or counterfactual distributions. The second part states that in absence of the observational distribution, no inferences can be made for higher layers of the PCH, leading to the following important corollary:
Corollary 2 (No Observational Distribution $\implies$ No Layer 2/3 Inference)
If the model’s distribution $\tilde{P}(V)$ differs from the true $P(V)$, no guarantees can be provided for the validity of the model’s interventional or counterfactual inferences.
The above corollary captures an important motivation for the benchmark constructed in this paper – even if we assume that an LLM has access to correct causal knowledge (a point we do not investigate in this manuscript), it still may not be able to perform causal inferences in case its observational distribution does not match the one in the real world. Interestingly, existing works in the literature treat language models as already possessing causal knowledge (de2025efficient), while our work provides evidence against such practice.
2 Benchmark Axes
This section is organized according to Fig. 2, and the different axes appearing therein. We first discuss (observational) probabilistic knowledge and its different dimensionalities (Sec. 2.1), followed by the description of datasets that are used (Sec. 2.2), and then the tasks constructed for eliciting models’ capabilities (Sec. 2.2.1). In Sec. 2.3, we discuss different models currently included in the benchmark.
2.1 Axis 1: Dimensionality of Probabilistic Knowledge
The observational distribution $P(V)$ over variables $V$ captures all probabilistic relationships among observed variables under passive observation. To test whether a model has access to the correct distribution, we may focus on different quantities, such as:
  1. (i)
    Marginal expectations: e.g., $\mathbb{E}[V_{1}]$, the average value of a single variable.
  2. (ii)
    Marginal distributions: e.g., $P(V_{1})$, the full distribution of a variable.
  3. (iii)
    Conditional expectations: e.g., $\mathbb{E}[V_{1}|V_{2}]$, the average value conditioned on another variable.
  4. (iv)
    Conditional distributions: e.g., $P(V_{1}|V_{2})$, the distribution of one variable given another.
  5. (v)
    Joint distributions: e.g., $P(V_{1},\dots,V_{k})$, the distribution of over the variables $V_{1},\dots,V_{k}$.
The above types of knowledge can be viewed as a hierarchy, with the level of required knowledge being progressively more refined. Our goal can be described as follows: we design tasks probing the above types of knowledge, based on large-scale datasets describing populations. For instance, we refer to a dataset $\mathcal{D}$, and consider the distribution over covariates $P(V)$ implied by $\mathcal{D}$. Then, we elicit responses from an LLM about the same distribution. The model’s distribution is labeled $\tilde{P}(V)$. Our goal is to then compare various aspects of $\tilde{P}(V)$ and $P(V)$, according to types of knowledge (i)-(v). In this way, we can systematically evaluate whether LLMs internalize such quantities over real-world populations.
2.2 Axis 2: Datasets Describing Populations
We leverage large-scale datasets that describe population level statistics in order to establish the ground truth observational distribution $P(V)$. After this, we probe the language model’s distribution $\tilde{P}(V)$ to draw comparisons. In this work, for the observational distributions $P(V)$, we make use of ten large, publicly available datasets that collectively describe diverse aspects of the US population, and are considered representative of national level statistics:
  1. (1)
    American Community Survey (ACS) 2023 (acs2023): Conducted by the US Census Bureau, the ACS collects detailed demographic, social, economic, and housing data annually. Our focus is on income, education, and employment information across demographics.
  2. (2)
    National Health and Nutrition Examination Survey (NHANES) 2021-2023 (nhanes2023): Administered by the Centers for Disease Control and Prevention (CDC), NHANES combines interviews and physical exams to assess health and nutrition of individuals in the US. We investigate obesity, diabetes, and dietary habits across demographics.
  3. (3)
    Behavioral Risk Factor Surveillance System (BRFSS) 2023 (brfss2023): A telephone survey system run by the CDC, tracking health-related risk behaviors and conditions. We investigate exercise habits, diabetes, blood pressure, asthma, cholesterol, visual/auditive impairments by US states.
  4. (4)
    Medical Expenditure Panel Survey (MEPS) 2023 (meps2023): A set of surveys conducted by the Agency for Healthcare Research and Quality (AHRQ), measuring health services use, expenditures, and insurance coverage. We investigate healthcare expenditure, utilization, and insurance across demographics.
  5. (5)
    National Survey on Drug Use and Health (NSDUH) 2023 (nsduh2023): Collected by Substance Abuse and Mental Health Services Administration (SAMHSA), the NSDUH provides information on substance use and mental health in the US. We investigate alcohol and drug use across demographics.
  6. (6)
    Survey of Consumer Finances (SCF) 2022 (scf2022): Sponsored by the Federal Reserve Board, the SCF provides detailed data on US household finances. We analyze food expenditure, home ownership, assets, and debt across demographics.
  7. (7)
    General Social Survey (GSS) 2022 (gss2022): Conducted by the National Opinion Research Center (NORC) at the University of Chicago, the GSS collects data on social attitudes, behaviors, and demographics of US adults. We investigate political views and party identification across age, sex, race, education, and income.
  8. (8)
    Department of Education (IPEDS) (ipeds2023): The Integrated Postsecondary Education Data System (IPEDS) collects data from colleges, universities, and technical schools. We investigate college degrees by sex.
  9. (9)
    Department of Labor BLS Data 2023 (bls2023): The Bureau of Labor Statistics of the US Department of Labor provides information about demographics across occupations. We investigate occupations by sex and race.
  10. (10)
    Federal Bureau of Investigation (FBI) Arrest Statistics (UCR Program): Compiled by the FBI’s Uniform Crime Reporting Program (UCR), this dataset contains arrest statistics across the US. We investigate crime rates by race and sex.
Each of the datasets either reports national-level statistics, or provides data on individuals. When individual data is available, sample weights are provided (determining how many persons in the overall US population the individual represents), allowing one to obtain representative results at the national level. Together, these datasets allow us to (approximately) establish population-level observational distributions across social, economic, educational, health, and behavioral domains. We next describe how different tasks in our benchmark are constructed.
2.2.1 Axis 2 Tasks: Responses and Scoring
Our benchmark consists of different tasks, and each task is associated with a dataset. The task is defined by the pair $V_{Y},V_{X}$, representing the conditional distribution $P(V_{Y}\mid V_{X})$ (here, $V_{X}$ could be possibly empty). Each task is also accompanied by a natural language prompt template $\pi$ (which takes a value of $V_{X}=v_{X}$ and returns a natural language question), and a set of valid answers in the domain of $V_{Y}$, labeled $\mathrm{dom}(V_{Y})$. For each task, we consider two ways of eliciting model responses, referred to as (a) question-answer (QA) prompting and (b) likelihood prompting.
Question & Answer Prompting.
The first approach for eliciting the model’s distribution is QA prompting, illustrated in the following example:
Example 2 (NSDUH: Marijuana Usage by Age)
The NSDUH dataset tracks various aspects of addiction and mental health for the US population. Suppose that $V_{Y}$ represents whether a person ever used marijuana, and $V_{X}$ represents age. In the benchmark, we are interested whether the LLM has access to the conditional distribution $P(V_{Y}\mid V_{X})$. To test this, the accompanying prompt template $\pi$ for the task is given by:
$\pi(v_{X})=\text{``Has a person aged \{$v_{X}$\} ever used marijuana?"}$ (1)
Once a specified value $v_{X}=16$ is chosen, this results in a prompt “Has a person aged 16 ever used marijuana?", which is given to the model. In the prompt, the model is provided with a set of possible answers, which in this case amounts to $\mathrm{dom}(V_{Y})=\{\text{no},\text{yes}\}$. For open-weights models, we can inspect the probabilities associated with each of the responses in the model’s next-token prediction. Alternatively, for closed models, the same question can be repeated multiple times, and we record all the provided answers, allowing us to reconstruct the distribution over answers. Finally, we compare the model’s distribution with the true $P(V_{Y}\mid V_{X})$ distribution from the NSDUH dataset.
Intuitively, the difficulty of the task should increase with larger dimensions of $V_{Y}$ and $V_{X}$. For instance, if $|V_{Y}|=1$, $|V_{X}|=0$, our task is to recover the 1-dimensional marginal distribution $P(V_{Y})$. If $|V_{Y}|=1,|V_{X}|>0$, we are trying to recover a 1-dimensional conditional distribution $P(V_{Y}\mid V_{X})$. The question templates $\pi$ are carefully designed, to reflect the original questions asked to individuals during data acquisition. In total, we constructed 75 tasks across 10 datasets for $|V_{Y}|=1,|V_{X}|=1$ (we refer to this as the low-dimensional setting), and 94 tasks across 4 datasets for $|V_{Y}|=1,|V_{X}|\in\{2,3,4,5\}$ and $V_{Y}$ binary (referred to as the high-dimensional setting). Fig. 2 (bottom left) offers some further examples, in addition to Ex. 2, while Appendix A contains the full list of tasks.
For QA prompting, our approach to obtain the model answers is to label each response in $\mathrm{dom}(V_{Y})$ with labels $A,B,C,\dots$, and then inspect the conditional probability associated with tokens $A,B,C,$ etc. Let $\mathcal{A}(v_{Y})$ denote the letter of the answer $v_{Y}$. The model’s probability associated with each answer $v_{Y}\in\mathrm{dom}(V_{Y})$ corresponding to label $\mathcal{A}({v}_{Y})$ in the prompt is then simply computed as
$\tilde{P}(V_{Y}=v_{Y}\mid V_{X}=v_{X})\overset{\Delta}{=}\frac{\tilde{P}^{\mathrm{nt}}(\text{answer }\mathcal{A}({v}_{Y}))}{\sum_{\tilde{v}_{Y}\in\mathrm{dom}(V_{Y})}\tilde{P}^{\mathrm{nt}}(\text{answer }\mathcal{A}(\tilde{v}_{Y}))},$ (2)
where $\tilde{P}^{\mathrm{nt}}$ is the model’s next-token probability function given the prompt.
Example 2 (continued – NSDUH: Marijuana Usage by Age)
Suppose that following the prompt
Has a person aged 16 years ever used marijuana? A. yes B. no
we find the model’s next-token probabilities to be $\tilde{P}^{\mathrm{nt}}(\text{answer A})=0.01,\tilde{P}^{\mathrm{nt}}(\text{answer B})=0.03$. Then, the model’s probability is computed as $\tilde{P}(V_{Y}=1\mid V_{X}=16)=\frac{0.01}{0.01+0.03}=25\%$.
This is a known strategy for eliciting model responses, used in numerous works on question answering (dominguez2024questioning; hendrycks2020measuring; santurkar2023whose), and many of the large benchmarks in the literature (hendrycks2020measuring; clark2018think; bisk2020piqa). As models may exhibit ordering bias (dominguez2024questioning), we consider all of the different $|\mathrm{dom}(V_{Y})|!$ permutations of labels $A,B,C,\dots$, over the answers $v_{Y}\in\mathrm{dom}(V_{Y})$, and average the probabilities accordingly (if the number of permutations $|\mathrm{dom}(V_{Y})|!>120$, we consider 120 random permutations). The above way of eliciting responses may be used for all models with open-source weights, which allows us to access next-token prediction probabilities. For the ground truth distribution $P$, in the low-dimensional setting we use a simple bin-counting estimator on the available dataset, whereas in the high-dimensional case we fit the distribution $P(V_{Y}\mid V_{X})$ using lightgbm (ke2017lightgbm), with actual values $P(V_{Y}=v_{Y}\mid V_{X}=v_{x})$ obtained out-of-sample, using 5-fold cross-validation.
Likelihood Prompting.
An alternative way of eliciting the distribution $P(V_{Y}=v_{Y}\mid V_{X}=v_{x})$ from the model is to inquire about the probability directly:
Example 2 (continued – NSDUH: Marijuana Usage by Age)
For the task of eliciting the distribution of marijuana usage by age group, we use the following prompt:
What is the probability that a person aged 16 years ever used marijuana?
    A. 0%   B. 0%-5%   ...   U. 95%-100%   V. 100%
We then inspect the next-token prediction probabilities in $\tilde{P}^{\mathrm{nt}}$, and choose the response with the highest probability:
$\mathcal{A}^{\text{resp}}=\operatorname*{arg\,max}_{\mathcal{A}\in\{A,\dots,V\}}\tilde{P}^{\mathrm{nt}}(\text{answer }\mathcal{A}).$ (3)
The chosen response letter $\mathcal{A}^{\text{resp}}$ is then mapped to either an interval, or a fixed value of 0% or 100%. For assigning the final predicted probability, for intervals, we take the midpoint. For instance, if the answer B. 0%-5% is chosen, we set the predicted probability to $0.025$.
For likelihood prompting, permutations over answer labels are not considered.
Scoring Strategy.
After explaining how model answers are obtained, we next develop a scoring strategy for each task based on the true distribution $P(V_{Y}\mid V_{X})$ and the elicited model’s distribution $\tilde{P}(V_{Y}\mid V_{X})$. For this purpose, we use the $L_{1}$-norm (instead of KL divergence, to avoid sensitivity to low-probability events and support mismatch (gibbs2002choosing)), and define the following distributional distance:
$D(\tilde{P}||P;V_{Y},V_{X})=\sum_{v_{X}}\sum_{v_{Y}}\Big|P(V_{Y}=v_{Y}\mid V_{X}=v_{X})-\tilde{P}(V_{Y}=v_{Y}\mid V_{X}=v_{X})\Big|P(V_{X}=v_{X})$ (4)
The above notion of distance allows the evaluation of how far $P,\tilde{P}$ are for the specific conditional $V_{Y}\mid V_{X}$. To obtain a normalized score, we compare the distance D(P||P~)D($|\tilde{P}$) to the distance of $P$ from a uniform distribution over the answers, labeled $P^{\text{unif}}$, defined as
$P^{\text{unif}}(V_{Y}=v_{Y}\mid V_{X}=v_{X})\overset{\Delta}{=}\frac{1}{|\mathrm{dom}(V_{Y})|}\;\forall v_{X},v_{Y}.$ (5)
[…]
Figure 3: Scoring strategy.
In this context, we refer to the $P^{\text{unif}}$ distribution as a random baseline. In the high-dimensional case, which focuses on binary $V_{Y}$, we add another baseline, namely a fixed 0/1 prediction depending on the dataset mean, given by $P^{0/1}(V_{Y}=1\mid V_{X}=v_{X})\overset{\Delta}{=}\mathbb{1}(\mathbbm{E}_{P}[V_{Y}]>0.5)$ for all $v_{Y},v_{X}$, where $\mathbbm{E}_{P}[V_{Y}]$ represent the true distribution mean. Therefore, the 0/1 baseline predicts a constant probability of 0 for all conditioning sets for outcomes with a marginal incidence $\mathbbm{E}_{P}[V_{Y}]\leq 0.5$, and predicts a constant $1$ whenever the marginal incidence is $\mathbbm{E}_{P}[V_{Y}]>0.5$. Such a baseline also corresponds to effectively no probabilistic knowledge. With these baselines in place, our scoring system is described next (illustrated in Fig. 3). As many datasets consist of samples of the population (and do not survey the entire population), there is some uncertainty on the ground truth distribution $P$. In the scoring, we account for this fact, and proceed as follows. We draw bootstrap samples of size $|\mathcal{D}|$ (size of the dataset) labeled $\mathcal{D}^{(b)}$, and extract the bootstrap ground truth distribution $P^{(b)}$. We then look at the distance $D(P||P^{(b)})$ of $P$ (based on the available data) and $P^{(b)}$ (obtained from bootstrapped data), across different samples (see green density in Fig. 3). The upper 5% quantile of the spread of $D(P||P^{(b)})$ is labeled with the perfect score $S=100$ – meaning that the score 100 is assigned if a model’s distribution cannot be statistically distinguished from the ground truth at the 5% significance level. For setting the score of 0, we use the minimum of the distances $D(P||P^{\text{unif}})$, $D(P||P^{0/1})$, with the latter only considered for binary outcomes. The model’s final score is then linearly interpolated between $S=0$ and S=100S={100}, as follows:
$S_{T}=100\times\max\Big(1-\frac{D(\tilde{P}\|P)}{\min{\{D(P^{\text{unif}}\|P),D(P^{\text{0,1}}\|P)\}}},0\Big).$ (6)
2.3 Axis 3: Models
Models we consider can be divided in two groups: open-weight models, whose trained parameters are publicly available, and closed-weight models, whose parameters remain proprietary and can be accessed only via an API. We focus mainly on open models, but also consider some closed models to ensure our benchmark covers a broad range of state-of-the-art models. While the benchmark evaluates zero-shot performance of models, a natural question is whether fine-tuning can help boost performance. Impact of fine-tuning is investigated in Appendix LABEL:appendix:fine-tune.
Open-weight Models.
Our open‐weight models include both the original pretrained checkpoints, and instruction‐tuned variants, which are further trained on human‐generated instruction–response pairs to improve their ability to follow instructions in user prompts. Particularly, we evaluate the following models: Mistral 7B (jiang2023mistral7b), LLaMA3 8B (grattafiori2024llama), LLaMA3 70B (grattafiori2024llama), Gemma3 27B (team2024gemma), DeepSeek 7B (bi2024deepseek), Phi-4 (abdin2024phi), and DeepSeek R1 32B (bi2024deepseek). In the main text, we focus on instruction-tuned models, while Appendix LABEL:appendix:base-vs-it compares instruction-tuned and pretrained models.
Closed‐weight Models.
Closed‐weight models are accessible only via API calls to proprietary weights. In this category, we evaluate state-of-the-art Reasoning Language Models (RLMs), including OpenAI’s o4-mini. We also include closed models with web-augmented capabilities, which rely on retrieval-augmented generation (RAG) (lewis2020retrieval), such as OpenAI’s GPT-4.1.
2.4 Benchmark Construction
Our benchmark is designed with reproducibility and extensibility as key principles. All dataset constructions are made available, while the modular structure of the benchmark allows easy addition of both new datasets and new models, facilitating broader future evaluations. Furthermore, model evaluation is streamlined: any model available through Hugging Face (wolf2019huggingface) can be benchmarked with a single line of code (see our HF repository), allowing for rapid inclusion of diverse architectures. At the time of submission, our benchmark consists of 10 pre-processed datasets, 169 tasks (75 low-dimensional and 94 high-dimensional), with 12 open-weight models evaluated.
3 Results
Overall Results.
Results summarizing the leaderboards for the low- and high-dimensional settings are shown in Fig. 4, and results can be viewed interactively on the project webpage. They indicate that the observational distributions $\tilde{P}(V)$ encoded in the LLMs are closer to the ground truth, real-world observational distributions $P(V)$ than the uniform distribution $P^{\text{unif}}(V)$ since models achieve scores above zero, suggesting some access to L1 knowledge. However, overall, the models’ performance is relatively poor. For QA prompting, best models scored an average of 22/100 points in the low-dimensional, and 17/100 in the high-dimensional setting, while for likelihood prompting, they score 41/100 and 33/100 respectively. Therefore, capabilities of LLMs for encoding real-world observational distributions seem limited, and using such knowledge for downstream tasks should be done with care.
Figure 4: Model leaderboard.
Furthermore, the analysis in Appendix LABEL:appendix:fine-tune shows that fine-tuning on samples from the ground truth dataset offers limited added benefit, raising interesting research questions on how to design approaches for improving the models’ L1 knowledge. In Appendix LABEL:appendix:fairness-impact we investigate the possible impact of debiasing techniques (tuning models for fairness and alignment), to understand if such methods may skew the results of our benchmark.
Low-Dimensional Setting & Performance by Dataset.
Another interesting aspect is the performance of models across different datasets (low-dimensional setting), visualized in Fig. 5. The best performance is observed over the FBI Arrest Statistics (sex/race distributions over crime types), ACS Census data (questions on employment, education, and income across demographics), and Departments of Education (sex by graduation degree) and Labor (sex/race by occupation). For three of these four datasets (FBI, BLS, IPEDS) the detailed statistics we queried are available on their websites, meaning that models could have had access to the exact probability tables used for constructing our questions. Appendix LABEL:appendix:data-contamination investigates whether improved performance may correspond to the actual data being available in the model’s training set, but we find no relation between question entropy (a proxy for whether a question was seen during pre-training (hendrycks2020measuring)) and model performance, implying that question memorization may not play an important role in context of our benchmark.
High-Dimensional Setting & Performance by Dimension.
We next inspect model performance with varying dimension in the high-dimensional setting (Fig. 6(a)). The pattern of the performances does not clearly indicate the curse of dimensionality, possibly due to the fact that each dimension $d$ has a relatively small number of tasks, and additional tasks may need to be added to investigate this effect in more detail. However, overall, models perform better in the low-dimensional setting than in the high-dimensional setting (Fig. 4).
Figure 5: Performance by dataset (low-dimensional).
Closed and Reasoning Models.
We next compare the performance of closed models vs. open-weight models. In particular, we evaluate two closed models from OpenAI – GPT-4.1 (achiam2023gpt) and o4-mini (openai2025o4mini), where the latter model is a reasoning language model (RLM) (besta2025reasoninglanguagemodelsblueprint). Additionally, we also include a version of DeepSeek R1 (guo2025deepseek) distilled using Qwen2.5 32B (hui2024qwen2). As the OpenAI models require access to an API, we evaluate them on a subset of the tasks in the high-dimensional setting (consisting of tasks with a univariate binary $V_{Y}$, and $|V_{X}|\in\{2,3,4,5\}$). In particular, we select 63 out of 94 tasks (the selected tasks are indicated in the Closed Eval column of Tab. LABEL:tab:high-dim), focusing on tasks for which at most 250 queries need to be evaluated on the model. To elicit the model’s distribution $\tilde{P}(V_{Y}\mid V_{X})$, we use likelihood-prompting, since closed models do not provide access to next-token prediction probabilities, making QA prompting possible only via Monte Carlo sampling, which is prohibitively expensive.  [^1] Monte Carlo sampling would increase the number of queries sent to the models and the associated cost by at least two orders of magnitude, due to the fact that even with $n_{mc}=100$ Monte Carlo samples, the 95% confidence interval for a probability $\tilde{P}(V_{Y}=1\mid V_{X}=v_{X})\in[0,1]$ is only guaranteed to be reduced to a width of approximately 0.1, still reflecting a high level of uncertainty.
(a) Performance by dimension.
(b) Performance of GPT-4.1 and o4-mini.
Figure 6: Model performance for (a) dimension; (b) closed models.
The performance of GPT-4.1, o4-mini, and DeepSeek R1 (Qwen2.5 32B distill) models on the 63 selected tasks is shown in Fig. 6(b). The figure shows that GPT-4.1 and o4-mini outperform the evaluated open-weights models. This may be explained by the fact that GPT-4.1 is the largest model we evaluated, in terms of the parameter count (even though the number of parameters is not officially published). Even though OpenAI’s models outperform the evaluated open-weights models, they still do not perform well, which we further contextualize in the discussion of baselines below.
Retrieval-Augmented Generation (RAG) (lewis2020retrieval).
Retrieval-augmented generation (RAG) is a technique in which an LLM is combined with an external retriever (lewis2020retrieval). In this case, the model does not have to rely only on its internal knowledge, but relevant documents can be fetched from a collection of texts. To investigate the impact of RAG on our benchmark, we used the GPT 4.1 model with web access (the model can search the web for each query). To test how RAG affects performance, we focused on 11 tasks on which the GPT 4.1 model without web-based RAG had a score of 0 (corresponding to task indices 40, 43, 57, 62, 63, 64, 65, 66, 67, 70, and 71 in Tab. LABEL:tab:high-dim). Here, the expectation is that RAG would help the model performance, and would result in a higher score. However, the results for the GPT 4.1 model with web-based RAG still showed a zero score for each task. This finding warrants further investigation of how RAG may improve the performance of models on our benchmark.
Baselines – how good are the models?
In the evaluation of our benchmark, two baselines are used, corresponding to the notion of no probabilistic knowledge. The first baseline was the uniform baseline $P^{\text{unif}}$, which is in the binary case given by:
$P^{\text{unif}}(V_{Y}=1\mid V_{X}=v_{X})=\frac{1}{2}\;\forall v_{X}.$ (7)
This baseline corresponds to random guessing. For high-dimensional tasks with binary outcomes, we also used the 0/1 baseline, which outputs either a probability 0 for events with a marginal probability $\leq 0.5$, or a probability $1$ for events with a marginal probability >0.5>0.5:
$P^{0/1}(V_{Y}=1\mid V_{X}=v_{X})=\mathbb{1}(\mathbbm{E}_{P}[V_{Y}]>0.5)\;\forall v_{X},$ (8)
where $\mathbbm{E}_{P}[V_{Y}]$ is the true mean of $V_{Y}$ (in the ground truth distribution $P$). In our scoring scheme, these baselines were used to determine the score $S=0$ for a given task (see Sec. 2.2.1). Clearly, the choice of baselines for determining the score $S=0$ has a strong effect on the evaluation. Here, another interesting baseline to consider is the mean-baseline, defined by:
$P^{\text{mean}}(V_{Y}=1\mid V_{X}=v_{X})=\mathbbm{E}_{P}[V_{Y}]\;\forall v_{X}.$ (9)
This baseline outputs a fixed value (the marginal mean of the outcome $V_{Y}$) for each $v_{X}\in\mathrm{dom}(V_{X})$. In words, this baseline is aware of the marginal probability of the event $V_{Y}$, but has no knowledge about the variation of $\mathbbm{E}[V_{Y}\mid V_{X}=v_{X}]$ according to $v_{X}$ around the marginal mean $\mathbbm{E}_{P}[V_{Y}]$. This baseline corresponds to some probabilistic knowledge (marginal mean), but without any knowledge over how the probability varies across population subsets. In Fig. 6(b), the performance of the mean-baseline with respect to the current evaluation is shown (orange horizontal line), indicating an average score of 46/100. This means that the inclusion of this baseline would effectively render the scores of all models investigated in this manuscript close to 0. This observation once again confirms that the current generation of LLMs do not possess robust knowledge on real-world observational distributions, regardless of the strategy (QA or likelihood prompting) for eliciting the model distributions $\tilde{P}(V_{Y}\mid V_{X})$.
Summary of Supplementary Analyses.
In addition to the main results reported above, several supplementary analyses (Appendices LABEL:appendix:base-vs-it-LABEL:appendix:data-contamination) further contextualize the findings. Appendix LABEL:appendix:base-vs-it compares base and instruction-tuned variants of the same models, showing that instruction tuning does not substantially change models’ performance, in some cases leading to worsening and in some to improvement. This suggests that instruction-tuning is not a key aspect when analyzing L1 knowledge. Appendix LABEL:appendix:fine-tune examines the effect of fine-tuning on samples from the ground-truth dataset and shows that such fine-tuning offers only limited benefit for improving models’ L1 knowledge. Appendix LABEL:appendix:fairness-impact evaluates the impact of debiasing and alignment techniques on benchmark performance, assessing whether such interventions may systematically skew the elicited distributions. Appendix LABEL:appendix:data-contamination investigates potential data contamination by relating question entropy to model performance on different tasks, and finds no evidence that memorization meaningfully contributes to performance. Together, these analyses support the overall conclusion that current LLMs possess limited knowledge of real-world observational distributions.
4 Conclusion
In this work, we introduced the first benchmark designed to systematically evaluate whether LLMs can provide knowledge on observational distributions across diverse domains; or, in other words, whether their statistics are aligned with the world they are intended to model. Our benchmark assesses language models on a range of associational quantities for which the ground truth can be relatively well established using large-scale population surveys. The benchmark tests the hypothesis that LLMs can serve as a proxy for real-world correlations. However, results across a suite of models suggest that models’ statistics are consistently and substantially misaligned with the ground truth. We hope that these findings motivate research developments to improve the statistical fidelity of current language models.
Implications on Causal Inferences.
As mentioned in the introduction, the recovery of observational distributions investigated in this work may be seen as related to questions of causal inference. Commonly, to adjust for confounding by a set of covariates $V_{Z}$, data analysts compute the propensity of a treatment $V_{X}$, that is, reweigh the samples with $P(V_{X}=v_{X}\mid V_{Z}=v_{Z})^{-1}$ to obtain estimates of a desired potential outcome (rubin1977assignment; rosenbaum1983central). Other strategies use a predictive model for the outcome for the outcome distribution $P(V_{Y}\mid V_{X},V_{Z})$ in order to allow for or improve causal inference (bang2005doubly). However, if such conditional distributions are not available to a model, eliciting them may be a futile exercise (see Cor. 2). Therefore, our work brings into question some recent approaches claiming to improve causal inferences using foundation models (de2025efficient). In our view, using AI models for causal inference may require a more sophisticated approach.
Limitations & Future Work.
We acknowledge some limitations of our work. Currently, a limited number of models are evaluated on our benchmark. However, by choosing a representative sample of models, we cover a range of models developed by different organizations. Furthermore, our emphasis on easy access and evaluation of models will facilitate the benchmarking of other models in future work. Another limitation is that high-dimensional inference is investigated across 4 datasets, and the number of contexts for evaluation should grow over time. Additionally, we mention that all datasets considered describe US populations, and future work should aim for wider geographic coverage. We further envision that our work will set the basis for evaluating capabilities of AI models for inference in Layers 2 & 3 of the PCH – to systematically test how observational distributions available to models affect downstream tasks such as causal or counterfactual inferences.
Supplementary Material for Epidemiology of Large Language Models: A Benchmark for Observational Distribution Knowledge
The source code for reproducing the benchmark can be found in our code repository https://github.com/dplecko/llm-epidemia. The repository includes a README file explaining all the steps for setting up the benchmark. Our Hugging Face repository https://huggingface.co/spaces/llm-observatory/llm-observatory-eval includes code for streamlined evaluations of models on the benchmark. Benchmark results can also be inspected on the project webpage https://llm-observatory.org/. For the Llama3 8B, Mistral 7B, and Phi4 models the experiments were run on a single NVIDIA H100 GPU. Evaluating both the low- and high-dimensional settings required <1 hour per model. Evaluating DeepSeek 7B (Chat version), Gemma3 27B, and LLama3 70B models was done on a node with four NVIDIA GH200 Grace Hopper chips, and running both settings required <1 hour of compute per model.
Appendix A Tasks
This appendix contains the full list of tasks considered in the benchmark, complementing the description of task construction from Sec. 2.2.1. In Tab. LABEL:tab:low-dim, the 75 tasks for the low-dimensional setting are listed, which are concerned with recovering $P(V_{Y}\mid V_{X})$ for univariate $V_{X},V_{Y}$. Tab. LABEL:tab:high-dim lists all the 94 tasks for the high-dimensional setting, recovering $P(V_{Y}\mid V_{X})$ for $V_{Y}$ univariate and binary, and $|V_{X}|\in\{2,3,4,5\}$.
Table 1: Low-dimensional tasks.
Task # Task Name
1 Census: Employment Status by Sex
2 Census: Employment Status by Race
3 Census: Employment Status by Age
4 Census: Employer by Sex
5 Census: Employer by Race
6 Census: Employer by Age
7 Census: Salary by Sex
8 Census: Salary by Race
9 Census: Salary by Age
10 Census: Education by Sex
11 Census: Education by Race
12 Census: Education by Age
13 BRFSS: Exercise by State
14 BRFSS: Diabetes by State
15 BRFSS: High BP by State
16 BRFSS: Asthma by State
17 BRFSS: Cholesterol by State
18 BRFSS: Visual Impairments by State
19 BRFSS: Hearing Impairments by State
20 BRFSS: Heart Attack by State
21 BRFSS: Stroke by State
22 Department of Education: Sex by Type of Degree
23 FBI Crime Statistics: Sex by Crime Type
24 FBI Crime Statistics: Race by Crime Type
25 GSS: Political View by Age
26 GSS: Political View by Race
27 GSS: Political View by Education
28 GSS: Political View by Income
29 GSS: Political View by Sex
30 GSS: Party Affiliation by Age
31 GSS: Party Affiliation by Race
32 GSS: Party Affiliation by Education
33 GSS: Party Affiliation by Income
34 GSS: Party Affiliation by Sex
35 Department of Labor: Sex by Occupation
36 Department of Labor: Race by Occupation
37 MEPS: Expenditure by Age Group
38 MEPS: Office-based Visits by Age Group
39 MEPS: Inpatient Visits by Age Group
40 MEPS: Dental Visits by Age Group
41 MEPS: Has Insurance by Age Group
42 MEPS: Expenditure by Race
43 MEPS: Office-based Visits by Race
44 MEPS: Inpatient Visits by Race
45 MEPS: Dental Visits by Race
46 MEPS: Has Insurance by Race
47 NHANES: Age by BMI Group
48 NHANES: Diabetes by BMI Group
49 NHANES: Diabetes by Age Group
50 NHANES: Weekly Alcohol Consumption by Age Group
51 NSDUH: Alcohol Use in Last Month by Age
52 NSDUH: Cigarette Use in Last Month by Age
53 NSDUH: Marijuana Ever Used by Age
54 NSDUH: Cocaine Ever Used by Age
55 NSDUH: Heroin Ever Used by Age
56 NSDUH: Alcohol Use in Last Month by Race
57 NSDUH: Cigarette Use in Last Month by Race
58 NSDUH: Marijuana Ever Used by Race
59 NSDUH: Cocaine Ever Used by Race
60 NSDUH: Heroin Ever Used by Race
61 SCF: Food Expenditure by Age Group
62 SCF: House Ownership by Age Group
63 SCF: Total Assets by Age Group
64 SCF: Debt by Age Group
65 SCF: Net Worth by Age Group
66 SCF: Food Expenditure by Race
67 SCF: House Ownership by Race
68 SCF: Total Assets by Race
69 SCF: Debt by Race
70 SCF: Net Worth by Race
71 SCF: Food Expenditure by Education
72 SCF: House Ownership by Education
73 SCF: Total Assets by Education
74 SCF: Debt by Education
75
Table 2: High-dimensional tasks. The third column indicates the task dimension $d=|V_{X}|$, and the fourth column indicates whether the task is used for the evaluation of closed models in Fig. 6(b).
Task # Task Name $d$ Closed Eval
1 BRFSS: Diabetes by Sex, Race 2
2 BRFSS: Diabetes by Age, Income 2
3 BRFSS: Diabetes by Age, Race 2
4 BRFSS: Diabetes by Sex, Income 2
5 BRFSS: Diabetes by Age, Sex, Race 3
6 BRFSS: Diabetes by Age, Race, Income 3
7 BRFSS: Diabetes by Age, Education, Income 3
8 BRFSS: Diabetes by Age, Sex, Income 3
9 BRFSS: Diabetes by Age, Education, Sex, Race 4
10 BRFSS: Diabetes by Age, Sex, Race, Income 4
11 BRFSS: Diabetes by Age, Education, Race, Income 4
12 BRFSS: Diabetes by Age, Education, Sex, Income 4
13 BRFSS: Diabetes by Age, Education, Sex, Race, Income 5
14 BRFSS: High Blood Pressure by Education, Race 2
15 BRFSS: High Blood Pressure by Age, Race 2
16 BRFSS: High Blood Pressure by Sex, Income 2
17 BRFSS: High Blood Pressure by Race, Income 2
18 BRFSS: High Blood Pressure by Age, Sex, Income 3
19 BRFSS: High Blood Pressure by Sex, Race, Income 3
20 BRFSS: High Blood Pressure by Age, Education, Race 3
21 BRFSS: High Blood Pressure by Education, Sex, Income 3
22 BRFSS: High Blood Pressure by Age, Sex, Race, Income 4
23 BRFSS: High Blood Pressure by Age, Education, Sex, Income 4
24 BRFSS: High Blood Pressure by Age, Education, Sex, Race 4
25 BRFSS: High Blood Pressure by Age, Education, Race, Income 4
26 BRFSS: High Blood Pressure by Age, Education, Sex, Race, Income 5
27 BRFSS: Depression by Race, Income 2
28 BRFSS: Depression by Sex, Race 2
29 BRFSS: Depression by Age, Income 2
30 BRFSS: Depression by Education, Race 2
31 BRFSS: Depression by Age, Race, Income 3
32 BRFSS: Depression by Education, Sex, Income 3
33 BRFSS: Depression by Age, Sex, Income 3
34 BRFSS: Depression by Sex, Race, Income 3
35 BRFSS: Depression by Age, Education, Sex, Race 4
36 BRFSS: Depression by Age, Education, Sex, Income 4
37 BRFSS: Depression by Education, Sex, Race, Income 4
38 BRFSS: Depression by Age, Sex, Race, Income 4
39 BRFSS: Depression by Age, Education, Sex, Race, Income 5
40 MEPS: Health Insurance by Education, Sex 2
41 MEPS: Health Insurance by Age, Sex 2
42 MEPS: Health Insurance by Age, Race 2
43 MEPS: Health Insurance by Education, Race 2
44 MEPS: Health Insurance by Age, Education 2
45 MEPS: Health Insurance by Sex, Race 2
46 MEPS: Health Insurance by Education, Sex, Race 3
47 MEPS: Health Insurance by Age, Sex, Race 3
48 MEPS: Health Insurance by Age, Education, Sex 3
49 MEPS: Health Insurance by Age, Education, Race 3
50 MEPS: Health Insurance by Age, Education, Sex, Race 4
51 NSDUH: Cigarette Use (Last 30d) by Education, Sex 2
52 NSDUH: Cigarette Use (Last 30d) by Age, Sex 2
53 NSDUH: Cigarette Use (Last 30d) by Age, Race 2
54 NSDUH: Cigarette Use (Last 30d) by Education, Race 2
55 NSDUH: Cigarette Use (Last 30d) by Age, Education 2
56 NSDUH: Cigarette Use (Last 30d) by Sex, Race 2
57 NSDUH: Cigarette Use (Last 30d) by Education, Sex, Race 3
58 NSDUH: Cigarette Use (Last 30d) by Age, Sex, Race 3
59 NSDUH: Cigarette Use (Last 30d) by Age, Education, Sex 3
60 NSDUH: Cigarette Use (Last 30d) by Age, Education, Race 3
61 NSDUH: Cigarette Use (Last 30d) by Age, Education, Sex, Race 4
62 NSDUH: Marijuana Use by Education, Sex 2
63 NSDUH: Marijuana Use by Age, Sex 2
64 NSDUH: Marijuana Use by Age, Race 2
65 NSDUH: Marijuana Use by Education, Race 2
66 NSDUH: Marijuana Use by Age, Education 2
67 NSDUH: Marijuana Use by Sex, Race 2
68 NSDUH: Marijuana Use by Age, Education, Race 3
69 NSDUH: Marijuana Use by Age, Sex, Race 3
70 NSDUH: Marijuana Use by Education, Sex, Race 3
71 NSDUH: Marijuana Use by Age, Education, Sex 3
72 NSDUH: Marijuana Use by Age, Education, Sex, Race 4
73 NSDUH: Cocaine Use by Education, Race 2
74 NSDUH: Cocaine Use by Education, Sex 2
75 NSDUH: Cocaine Use by Age, Race 2
76 NSDUH: Cocaine Use by Age, Education 2
77 NSDUH: Cocaine Use by Sex, Race 2
78 NSDUH: Cocaine Use by Age, Sex 2
79 NSDUH: Cocaine Use by Age, Sex, Race 3
80 NSDUH: Cocaine Use by Age, Education, Race 3
81 NSDUH: Cocaine Use by Age, Education, Sex 3
82 NSDUH: Cocaine Use by Education, Sex, Race 3
83 NSDUH: Cocaine Use by Age, Education, Sex, Race 4
84 SCF: Home Ownership by Education, Sex 2
85 SCF: Home Ownership by Age, Sex 2
86 SCF: Home Ownership by Age, Race 2
87 SCF: Home Ownership by Education, Race 2
88 SCF: Home Ownership by Age, Education 2
89 SCF: Home Ownership by Sex, Race 2
90 SCF: Home Ownership by Education, Sex, Race 3
91 SCF: Home Ownership by Age, Sex, Race 3
92 SCF: Home Ownership by Age, Education, Sex 3
93 SCF: Home Ownership by Age, Education, Race 3
94 SCF: Home Ownership by Age, Education, Sex, Race 4
Appendix B Comparison of Base & Instruction-Tuned Models
In this appendix, we compare the performance of base models vs. instruction-tuned models optimized via either self-instruction (wang2022self) or supervised fine-tuning (ouyang2022training) followed by reinforcement learning with human feedback (RLHF) (christiano2017deep). In the main text, we analyzed the performance of instruction-tuned models on our benchmark: Mistral 7B (jiang2023mistral7b), LLaMA3 8B (grattafiori2024llama), LLaMA3 70B (grattafiori2024llama), Gemma3 27B (team2025gemma), DeepSeek 7B (bi2024deepseek), and Phi-4 (abdin2024phi). For each of these models, we compare the model’s performance against the corresponding base (pre-trained) model. The only exception is the Phi-4 model, for which a base model is not released due to safety concerns (abdin2024phi).
For evaluating the models on our benchmark, we focus on the 94 tasks in the high-dimensional setting (see Sec. 2.2.1 and Tab. LABEL:tab:high-dim for details). The results of comparing base vs. instruction-tuned models are shown in Fig. LABEL:fig:base-vs-it, which again indicates poor performance across all models. We note that, for all families of models (here, a family refers to a pair of base and instruction-tuned models), instruction-tuned models perform equal or better than base models. Furthermore, a comparison of base model’s vs. instruction-tuned model’s scores for each task is shown in Fig. LABEL:fig:base-vs-it-bytask. The figure indicates that for a number of tasks, instruction-tuning improves performance from a zero score to a non-zero score. Less commonly, instruction tuning makes performance worse (some instances are observed for DeepSeek and Gemma 3 model families).
Appendix C Does Fine-Tuning Improve Performance?
In this appendix, we evaluate whether fine-tuning helps the performance of LLMs on our benchmark. Given the relatively poor performance of models across a range of tasks, a natural question arises whether additional training could improve models’ knowledge of observational distributions. We investigate this hypothesis by fine-tuning a LLaMA3 8B and Mistral 7B instruction-tuned models on synthetic data generated from our ground truth datasets, and comparing the performance to the baseline pretrained models.
C.1 Synthetic Data Creation
Our first step is to create a synthetic dataset that simulates the type of text a model might be given access to during its pretraining stage. To generate such that, we use a pretrained Gemma3 27B instruction model (team2025gemma) and focus on the NSDUH and BRFSS data. Specifically, we sample different individuals from the dataset, reflecting the observational distribution. Upon selecting an individual, we generate a text paragraph representing the person and their associated covariates.
C.1.1 NSDUH Data
For NSDUH data, we use the following prompt template for querying the Gemma3 27B model:
Prompt
You are a data generator. Follow the RULES strictly:
1) Write a single narrative enclosed in <story>...</story>.
2) Do NOT include headings, lists, analysis, or
any text outside the tags.
3) Mention ALL facts given below exactly once (age, sex, race,
   education, cigarette and alcohol last-month use,
   marijuana, cocaine, and heroin ever use).
4) Keep it under 200 words.
FACTS:
- age: {age}
- sex: {sex}
- race: {race}
- education: {edu}
- alcohol last month: {alc_monthly}
- cigarettes last month: {cig_monthly}
- alcohol last month: {cig_monthly}
- marijuana ever: {mj_ever}
- cocaine ever: {coc_ever}
- heroin ever: {her_ever}
OUTPUT FORMAT:
<story> (your narrative here) </story>
Such a prompt is used since tasks look at cigarette, alcohol, marijuana, cocaine, and heroin usage across demographics and therefore each generated textual note must contain all of this information. In particular, We create a dataset of 2,500 samples. An example of a generated data sample is shown below, showcasing the type of data to be used for fine-tuning:
Example Output Ethan, a 22-year-old White male, was navigating the complexities of early adulthood. He’d started some college but hadn’t yet earned a degree, feeling unsure about his path. He worked part-time at a local bookstore, enjoying the quiet atmosphere and the smell of old paper. Ethan wasn’t a smoker or a drinker; he hadn’t used cigarettes or alcohol in the last month. He’d never experimented with marijuana, cocaine, or heroin, preferring to stay clear of substances. His friends were often surprised by this, but Ethan was content with his choices. He spent his free time sketching in his notebook, dreaming of becoming an illustrator, and occasionally volunteering at the animal shelter.
C.1.2 BRFSS Data
For the BRFSS data, we use a very similar prompt template for querying the Gemma3 27B model, but change the facts generated by the model:
Prompt
FACTS:
- state: {state}
- age: {age}
- sex: {sex}
- race: {race}
- education: {education}
- income: {income}
- smoking status: {smoker}
- BMI: {bmi}
- exercise monthly: {exercise_monthly}
- poor mental health: {poor_mental_health}
- diabetes: {diabetes}
- high blood pressure: {high_bp}
- asthma: {asthma}
- cholesterol: {cholesterol}
- heart attack: {heart_attack}
- stroke: {stroke}
- depression: {depression}
Since different tasks look at health status across demographics, so the textual data must include all of these attributes. Again, we create a dataset of 2,500 samples, such as:
Example Output Eleanor, a 52.0 year old White female, lived a relatively quiet life in Ohio. She was a college graduate earning between \$50–100k annually. She had never been diagnosed with diabetes, high blood pressure, asthma, or high cholesterol. Thankfully, she’d also avoided both a heart attack and stroke. Eleanor did not smoke and maintained a BMI of 25.84, though she admitted her exercise routine was limited to about monthly activity. She reported good physical health and fortunately, had no history of depression. Her mental health was stable; she didn’t suffer from poor mental health.
C.2 Fine-Tuning
By fine-tuning a decoder-only model (LLama3 8B or Mistral 7B) on the above two datasets, we test whether a model can internalize observational knowledge by learning to do causal language modeling on narrative text. An alternative fine-tuning approach would be to update model weights directly on summarized statistics. However, this way of training would no longer represent embedding probabilistic knowledge (which our benchmark aims to assess) into the model, and would instead embed factual knowledge.
Standard supervised finetuning (SFT) is used for both LLama3 8B and Mistral 7B, which updates all model weights. The cross-entropy loss for next token prediction is optimized. Each model is trained for 6 epochs using a learning rate of $5\cdot 10^{-5}$ and the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$. We apply a linear learning rate scheduler and monitor the loss on both training and validation splits. We save the model checkpoints that achieve the lowest evaluation loss. Loss statistics over the 6 epochs are shown in Fig. LABEL:fig:loss-history.
As the loss behavior indicates, the fine-tuning approach reduces both training and evaluation losses early on, after which overfitting begins, and the training loss continues to decrease while evaluation loss starts to increase. This observation is consistent across models and datasets, showing adequate fine-tuning behavior. Model checkpoints saved using early stopping are indicated with a vertical red line.
C.3 Evaluation & Observations
In the last step, we evaluate the fine-tuned models on all of the NSDUH and BRFSS tasks, and compare the task performances to those of pretrained LLama3 8B and Mistral 7B models. The results are shown in Fig. LABEL:fig:ft-results and Fig. LABEL:fig:ft-results-brfss for both settings (low- and high-dimensional) and prompting techniques (question-answer and likelihood prompting) for the respective datasets.
C.3.1 NSDUH Data
For QA prompting, fine-tuning does not seem to improve performance across either low- or high-dimensional settings – very few and modest score increases are recorded for both models. At the same time, in the high-dimensional setting, for some tasks where a pretrained LLama3 8B model scored above 0 the performance drops with fine-tuning.
For likelihood prompting, the impact of fine-tuning is different for low-dimensional and high-dimensional settings. In the low-dimensional setting, FT increases the performance on 6 out of 10 tasks for LLama3 8B and 3 out of 10 for Mistral 7B. In the high-dimensional setting, the performance is not improved on any of the 33 tasks for either model, so both the pretrained and fine-tuned models score exactly 0 on each task. Therefore, some FT gains are observed for likelihood prompting in the low-dimensional setting.
C.3.2 BRFSS Data
For QA prompting, fine-tuning leads to some performance improvements (particularly for Mistral 7B, see Fig. LABEL:fig:ft-mistral-brfss), but these do not appear consistent across tasks. Furthermore, there are also instances of a drop in performance in certain tasks. For Llama3 8B the results are also not conclusive, with some improvements on tasks where the pretrained model score 0, but also a drastic drop in performance across tasks where the pretrained model scored more favorably.
For likelihood prompting, the impact of fine-tuning is again different for low-dimensional and high-dimensional settings. In the low-dimensional setting, FT increases the performance on 4 out of 10 tasks for LLama3 8B, yet it also reduces the performance to zero on one task. Similarly, for Mistral 7B FT reduces the performance on 3 out of 10 tasks, and show an increase for one task. In the high-dimensional setting, the performance is not improved on any of the 33 tasks for either model; in fact the fine-tuned Mistral 7B model performs worse than the pretrained one, and has a consistent 0 score across all tasks. Thus, no consistent improvements are observed across different settings.
C.3.3 Fine-Tuning Takeaways
One would perhaps hypothesize that fine-tuning on the task of causal language modeling would have a greater impact and improve upon performance for the QA prompting, since this prompting technique more directly elicits the model’s internal probability from next token predictions (instead of asking for a likelihood), and the fine-tuning data encodes the correct underlying distribution in this form. However, the observed results do not support this hypothesis. Therefore, our investigation in this appendix offers preliminary evidence that fine-tuning models on text generated from the correct observational distribution does not ensure that models internalize knowledge of these distributions. Future work is needed to extend these findings to other models and data sources, and to investigate methods that may improve the abilities of models in terms of probabilistic L1 knowledge.
Appendix D Fairness Impacts
In this appendix, we assess whether the presence of protected attributes in a task influences model performance. Language models are often trained with a variety of safety-oriented or debiasing interventions, e.g., safety tuning, data mixture filtering, or RLHF reward shaping (bolukbasi2016man; zhao2018learning; glaese2022improving). These techniques may cause models to behave differently when a prompt mentions sensitive attributes such as sex or race. If such mechanisms are active, they could in principle deflate scores on tasks where protected attributes appear directly in the question; hence it is important to verify that the main results are not driven by this effect. To examine this, we partition the benchmark tasks into two groups: (i) tasks that reference attributes sex or race (118/169 tasks), and (ii) tasks that contain no such attributes (51/169 tasks). For each model and prompting mode, we compute average performance separately in these two groups.
The results are shown in Fig. LABEL:fig:fairness-impact, with a split in sensitive and non-sensitive tasks. Overall, most models perform slightly better on non-sensitive tasks, although not all. Furthermore, the performance difference appears to be modest, and the overall pattern of findings remains unchanged. These results support the fact that the benchmark’s findings are not driven purely by the debiasing techniques applied on protected attributes. Tasks involving sex or race exhibit slightly worse scores, but the qualitative picture of model underperformance persists across both categories.
Appendix E Data Contamination
In this appendix, we analyze potential data contamination during pretraining, that is, whether during pretraining models may have seen the exact questions used in the benchmark. There is a non-zero chance for this, as language models are trained on massive text corpora. As discussed in (hendrycks2020measuring), if a model simply recalls a memorized question-answer pair, this would be reflected in lower question entropy and a higher score. Such data contamination would also make reported model performance appear more favorable than the true underlying model capability.
In light of this, we investigate whether there is a relationship between question entropy and task performance. We focus on likelihood prompting, as this is the more likely question format that models may have encountered during training. In Fig. LABEL:fig:data-contamination, we plot the score and question entropy for each low-dimensional and high-dimensional task for LLaMA3 8B and Phi-4 models. As the figure illustrates, there is an absence of a positive correlation between score and question entropy, suggesting that there is no evidence that exact question–answer pairs were memorized during pretraining. For a further discussion on investigating question memorization using entropy, we refer the reader to (hendrycks2020measuring) and (brown2020language).